# Concentration Inequalities for Branching Random Walks

### v5

@@ 3 Concentration inequalities in the context of BRW
i.e. $E^{(M)}_{i}(X_{\ell})$ is Lipschitz continuous function of $u_{i}$.
Recall that in the proof of Theorem 2.2, smooth $\mu_{i}(u_{i})$ and $\nu_{i}(u_{i})$ are employed to prove “uniqueness of solution” (in RW “local” smoothness is “global” smoothness because it won’t change with generations—steps). In the case of BRW, we need smoothness of $\mu_{i}^{\,\!{}^{(M)}}(u_{i}):=E_{i}^{\,\!{}^{(M)}}(X_{i+1}|u_{i})$, $\nu_{i}^{\,\!{}_{(M)}}(u_{i}):=E_{i}^{\,\!{}^{(M)}}(X_{i+1}^{2}|u_{i})$, and of $\mathbb{E}_{i}^{(M)}\!\big(m(u_{i+M})\mid u_{i}\big)$. To this end, we first transform the “local” smoothness (given as hypotheses in the theorem) into “global” smoothness; i.e. to prove $\mu_{i}^{\,\!{}^{(M)}}(u_{i})$, $\nu_{i}^{\!{}^{(M)}}(u_{i})$ and $\beta_{i}^{\,\!{}^{(M)}}(u_{i})$ are smooth functions of $u_{i}$. We proceed by induction on the generations, recalling that the hypotheses in the theorem, the assumptions on $\beta_{i}$, $\mu_{i}$ and $\nu_{i}$ (without “(M)”), ensure that the base case (generation 0) holds. The following concerns generation $M$, where we use $n$ instead of $i$ for easy of comparison with the proof of Theorem 2.2.

### v6

@@ 3 Concentration inequalities in the context of BRW
i.e. $\mathbb{E}^{(M)}_{i}(X_{\ell})$ is Lipschitz continuous function of $u_{i}$.
Recall that in the proof of Theorem 2.2, smooth $\mu_{i}(u_{i})$ and $\nu_{i}(u_{i})$ are employed to prove “uniqueness of solution” (in RW “local” smoothness is “global” smoothness because it won’t change with generations—steps). In the case of BRW, we need smoothness of $\mu_{i}^{\,\!{}^{(M)}}(u_{i}):=\mathbb{E}_{i}^{\,\!{}^{(M)}}(X_{i+1}|u_{i})$, $\nu_{i}^{\,\!{}_{(M)}}(u_{i}):=\mathbb{E}_{i}^{\,\!{}^{(M)}}(X_{i+1}^{2}|u_{i})$, and of $\mathbb{E}_{i}^{(M)}\!\big(m(u_{i+M})\mid u_{i}\big)$. To this end, we first transform the “local” smoothness (given as hypotheses in the theorem) into “global” smoothness; i.e. to prove $\mu_{i}^{\,\!{}^{(M)}}(u_{i})$, $\nu_{i}^{\!{}^{(M)}}(u_{i})$ and $\beta_{i}^{\,\!{}^{(M)}}(u_{i})$ are smooth functions of $u_{i}$. We proceed by induction on the generations, recalling that the hypotheses in the theorem, the assumptions on $\beta_{i}$, $\mu_{i}$ and $\nu_{i}$ (without “(M)”), ensure that the base case (generation 0) holds. The following concerns generation $M$, where we use $n$ instead of $i$ for easy of comparison with the proof of Theorem 2.2.